ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 108, 1998

Auinger, K. and M. B. Szendrei E-solid e-varieties are not finitely based

327

Benjamini, I. Expanders are not hyperbolic  33

Blume, F. Minimal rates of entropy convergence for completely ergodic systems

1

Bothe, D. Multivalued perturbations of m-accretive differential inclusions

109

Bruin, H. and J. Hawkins Examples of expanding C^{1} maps having no σ-finite invariant measure equivalent to Lebesgue

83

Eigen, S., A. Hajian and K. Halverson Multiple recurrence and infinite measure preserving odometers

37

Farmer, D. and D. Lieman The GL(3) Mellin transform for twisted non-cuspidal forms of higher level

291

Finet, C. and R. Payá Norm attaining operators from L_{1} into L_{∞}  139

Fonf, V. P. and J. Lindenstrauss Some results on infinitedimensional convexity

13

Foreman, M. An ℵ_{1}-dense ideal on ℵ_{2}  253
Fulman, I. and P. S. Muhly Bimodules, spectra, and Fell bundles  193

Hajian, A. See S. Eigen

Halverson, K. See S. Eigen

Hawkins, J. See H. Bruin

Judd, R. and E. Odell Concerning Bourgain’s ℓ_{1}-index of a Banach space

145

Lieman, D. See D. Farmer

Lindenstrauss, J. See V. P. Fonf

Muhly, P. S. See I. Fulman

Novik, I. Upper bound theorems for homology manifolds  45

Odell, E. See R. Judd

Payá, R. See C. Finet

Quine, J. R. and P. L. Zhang Extremal symplectic lattices  237

Saksman, E. and H.-O. Tylli Rigidity of commutators and elementary operators on Calkin algebras 217 Szendrei, M. B. See K. Auinger

Talagrand, M. Selecting a proportion of characters  173

Terenzi, P. A complement to the paper “A positive answer to the basis problem”

345

Tylli, H.-O. See E. Saksman Zhang, P. L. See J. R. Quine